OpenStudy (anonymous):

A certain mountain has an elevation of 19,257 feet. In 1918 the glacier of the peak covered 8 acres. By 2001 the glacier had melted to only one acre. (a) Assume that this glacier melted at a constant rate each year. Find this yearly rate. (b) Use your answer from part (a) to write a linear equation that gives the average A of this glacier T years past 1918.

OpenStudy (amistre64):

difference in acres, divided by the difference in years; gives an average yearly rate

OpenStudy (anonymous):

so it would be 8-1=7 7 divided by 2001-1918=7/83 am I correct

OpenStudy (amistre64):

1 - 8 ----------- = -7/83 2001 - 1918 its decreasing at a rate of 7/83 per year, yes

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

youre welcome